The sum of two numbers is fifteen. One less than three times the smaller is equal to the larger. Find the two numbers.
Found 2 solutions by vleith, stanbon:
Answer by vleith(2983)
Answer by stanbon(75887)

The sum of two numbers is fifteen. One less than three times the smaller is equal to the larger. Find the two numbers.
-------------------------------------------------
Equations:
x + y = 15
3x - 1 = y
----------------
Substitute 3x-1 in place of "y" to solve for "x":
x + 3x-1 = 15
4x = 16
x = 4
--------------
Substitute to solve for "Y":
y = 3*4-1 = 11
==============================
Solution: (4,11)
==============================
Cheers,
Stan H.